Use the following data set to solve for the level of GDP (Y), the level savings (S), the level of investment (I), and the level of government purchase of goods and services (G). YD= 12,000 TA =2,000 TR= 0 C= 7,000 G= (Y)*.2 NX =400  


1
Expert's answer
2021-04-09T06:59:04-0400

 GDP = C + I + G + NX

G=government spending

I = GDP − C − G − NX ).

Government Saving (Sgovt)  = T – TR- INT – G

YD = disposable income. S = savings. YD = Y + TR – TA. Y = national income. TR = government transfer payments.

(a) Level of GDP (Y)

YD = Y + TR – TA.

YD=12000

TR=0

TA=2000

"12000=Y+0-2000"

"Y=12000+2000=14000"

(b)Level of Savings

 =T − G − TR

"2000-28000-0=-26000"

(c)The level of investment

GDP = C + I + G + NX, thus,

I=Y-C-G-NX

I="14000-7000-28000-400=-21400"


(d)The level of government purchase of goods and services (G)

G=Y×2

Y=14000

G="14000\u00d72=28000"